BexleyBrooks:

Need some help on Math. 1. -3.5y-6.2y=87.3 What is D? 2. -1/4 d - -2/5d=39 What is D? 3. -9.76d - (-12.81d)=8.54 What is D?

BexleyBrooks:

Its combining like terms I think is what its called

luhivqqcherry:

-3.5y - 6.2y = - 87.3 or, -9.7y = -87.3 or, y = -87.3/-9.7 or, y = 9.

luhivqqcherry:

2. -1/4 d - -2/5d=39 What is D? D is = 60

luhivqqcherry:

3. -9.76d - (-12.81d)=8.54 What is D? for number 3 D = 2.8